# MMA-Diffusion: MultiModal Attack on Diffusion Models

### v3

@@ Abstract
In recent years, Text-to-Image (T2I) models have seen remarkable advancements, gaining widespread adoption. However, this progress has inadvertently opened avenues for potential misuse, particularly in generating inappropriate or Not-Safe-For-Work (NSFW) content. Our work introduces MMA-Diffusion, a framework that presents a significant and realistic threat to the security of T2I models by effectively circumventing current defensive measures in both open-source models and commercial online services. Unlike previous approaches, MMA-Diffusion leverages both textual and visual modalities to bypass safeguards like prompt filters and post-hoc safety checkers, thus exposing and highlighting the vulnerabilities in existing defense mechanisms.
Machine Learning
[…]
Warning: This paper includes AI-generated content that has the potential to be harmful. It is important to exercise caution and approach the content with sensitivity. The purpose of including these elements is solely for the purpose of research and academic discussion. Please be aware that these generated materials may contain explicit, offensive, or controversial content. Reader discretion is advised.
@@ 1 Introduction
In the rapidly evolving landscape of text-to-image (T2I) generation, diffusion models such as Stable Diffusion (SD) (Rombach et al., 2022) and Midjounery (Midjourney, 2023) have marked a paradigm shift. These models have revolutionized digital creativity by generating strikingly realistic images, yet they also pose significant security challenges. Notably, the potential misuse of these models for generating Not-Safe-For-Work (NSFW) contents (Schramowski et al., 2023; Qu et al., 2023; Saharia et al., 2022), such as adult materials, violence, and politically sensitive imagery, is a serious concern.
In response to these concerns, developers of T2I models have implemented preventive measures like prompt filters and post-synthesis safety checks. While effective to an extent, the resilience of these measures against sophisticated adversarial attacks remains a topic of intense debate and investigation. Our study delves into this pressing issue by introducing MMA-Diffusion, a framework designed to rigorously test and challenge the security of T2I models. Unlike conventional methods that make subtle prompt modifications (Liu et al., 2023a; Gao et al., 2023; Zhuang et al., 2023; Kou et al., 2023), MMA-Diffusion adopts a systematic attack approach. It enables users to generate unrestricted adversarial prompts and craft image perturbations, thereby circumventing existing safety protocols.
The technical prowess of MMA-Diffusion lies in its dual-modal attack strategy. We develop an advanced text modality attack mechanism that intricately alters textual prompts while maintaining their semantic intent, allowing for the generation of targeted NSFW content without being flagged by existing filters, as demonstrated in  Figure 1(a). On the image modality front, MMA-Diffusion utilizes a novel perturbation technique that subtly alters image characteristics in a manner undetectable to the human eye but significant enough to bypass post-processing safety algorithms, as illustrated in  Figure 1(b).
@@
Overall, the contributions of this work include: 1 We present a novel multimodal systematic attack that effectively bypasses prompt filters and safety checkers, highlighting a significant security issue in T2I models. 2 In the textual modality, we craft an adversarial prompt generation method that can deceive the prompt filter while remaining semantically similar to the target. For the image modality, we devise an attack that proficiently bypasses the post-hoc defense mechanism. 3 We evaluate various T2I models, encompassing popular open-source models and online platforms and demonstrate the effectiveness of the proposed MMA-Diffusion. For example, 10-query black-box attack can achieve a 83.33% and 90% success rate w.r.t Midjounery (Midjourney, 2023) and Leonardo.Ai (Leonardo.Ai, 2023).
@@ 2 Related Work
Adversarial attacks on T2I models. To the best of our knowledge, current research does not extensively explore attacks in the image modality for NSFW content generation with T2I models. Most existing studies on adversarial attacks in T2I models, such as (Qu et al., 2023; Gao et al., 2023; Zhuang et al., 2023; Kou et al., 2023; Liu et al., 2023b), have predominantly focused on text modification to probe functional vulnerabilities. These explorations encompass impacts from diminishing synthetic quality (Liu et al., 2023b; Zhang et al., 2023a) to distorting or eliminating objects (Zhuang et al., 2023; Liu et al., 2023b), and impairing image fidelity (Liu et al., 2023a). However, they do not target generating NSFW-specific materials like pornography, violence, politics, racism, or horror. Recent works such as UnlearnDiff (Zhang et al., 2023b) and Ring-A-Bell (Tsai et al., 2023) have started to consider the misuse of T2I models for generating NSFW content. UnlearnDiff primarily examines concept-erased diffusion models (Rombach et al., 2022; Gandikota et al., 2023; Kumari et al., 2023; Schramowski et al., 2023) and does not extend to other defense strategies. Conversely, Ring-A-Bell explores inducing T2I models to generate NSFW concepts but lacks precision in controlling the details of the synthesis. However, none of them considers attacks that can bypass both the prompt filter and the post-hoc safety mechanisms while still producing high-quality NSFW content tailored to specific semantic prompts. This paper demonstrates the feasibility of such attacks, highlighting their general applicability across a variety of T2I models.
@@
Defensive methods. Various T2I models implement distinct countermeasures to mitigate user abuse. Notably, popular online T2I services like Midjourney (Midjourney, 2023) and Leonardo.Ai (Leonardo.Ai, 2023) employ AI moderators to screen potentially harmful prompts. This proactive approach targets the prevention of NSFW content generation at the input stage. Another defensive strategy involves post-hoc safety checkers, exemplified by the one integrated into Stable Diffusion (SD) (Safety-checker, 2023; Rando et al., 2022a). Unlike AI moderators, these checkers function at the output stage, scrutinizing generated images to detect and obfuscate NSFW elements. Additionally, some novel mitigation methods lie in the concept-erased diffusion (Gandikota et al., 2023; Kumari et al., 2023; Schramowski et al., 2023). These methods differ fundamentally from external safety measures as they modify the model’s inference guidance or utilize fine-tuning to actively suppress NSFW concepts. However, they may not entirely eliminate NSFW content and could inadvertently affect the quality of benign images (Zhang et al., 2023b; Lee et al., 2023; Schramowski et al., 2023). This paper presents a multimodal attack that breaches both prompt filters and post-hoc safety checkers, which is also applicable to concept-erased diffusion models (e.g., SLD (Schramowski et al., 2023)), exposing the risk of T2I models and related online services.
Figure 2: Overview of the proposed framework. T2I models incorporate safety mechanisms, including (a) prompt filters to prohibit unsafe prompts/words, e.g.naked,” and (b) post-hoc safety checkers to prevent explicit synthesis. (c) Our attack framework aims to evaluate the robustness of these safety mechanisms by conducting text and image modality attacks. Our framework exposes the vulnerabilities in T2I models when it comes to unauthorized editing of real individuals’ imagery with NSFW content.
@@ 3.1 Threat Model
  • White-Box Settings: Here, adversaries utilize open-source T2I models like SDv1.5 (SDv1.5, 2023) for image generation. With full access to the model’s architecture and checkpoint, attackers can conduct in-depth explorations and manipulations for sophisticated attacks.
@@ 3.2 Approach Overview
In this paper, we focus on the attack that enables T2I models to generate high-quality NSFW content, thereby exposing the potential misuse risks of them, as in  Figure 2. Specifically, we assume that the attacker describes the content they wish to generate through plain text. The attack is considered successful only if the model generates NSFW content that aligns with the description.
To make the attack more realistic, we assume that the T2I model or the online service adopts two defense methods, namely: prompt filter, as in  Figure 2 (a) and post-hoc safety checker, as in  Figure 2 (b). For situations where only the prompt filter is present, such as (Leonardo.Ai, 2023), we employ a text-modal attack. For situations where only the post-hoc safety checker is present, such as SD (Rombach et al., 2022), we utilize an image-modal attack. For models that adopt both modalities of defense, we can simultaneously use both attack methods to achieve a stronger effect, as in  Figure 2 (c).
@@
3.3 Text-Modal Attack
T2I models typically rely on a pre-trained text encoder, $\tau_{\theta}(\cdot)$, to transforms natural language input $\mathbf{p}$ into a latent vector, denoted as $\tau_{\theta}(\mathbf{p})\in\mathbb{R}^{d}$, which is key to determining the semantics of the image synthesis (Nichol et al., 2022). The input sequence is $\mathbf{p}=[p_{1},p_{2},...,p_{L}]\in\mathbb{N}^{L}$, where $p_{i}\in\{0,1,...,|V|-1\}$ is the $i^{\text{th}}$ token’s index, $V$ is the vocabulary codebook, $|V|$ is the vocabulary size, and $L$ is the prompt length. This mapping from $\mathbb{N}^{L}$ to $\mathbb{R}^{d}$ provides a large search space for an attack, given a sufficiently large vocabulary pool $V$ and no additional constraints on $\mathbf{p}_{\text{adv}}$, therefore enabling free-style adversarial prompt manipulation.
The target of the text-modal attack is to evade the prompt filter while keeping the functionality guiding the T2I model for the desired NSFW content. Specifically, we set this original NSFW prompt as the target prompt, denoted as $\mathbf{p}_{\text{tar}}$ (e.g., “A completely naked Trump stands on the grass”). MMA-Diffusion assumes the prompt filter is implemented by filtering the prompts according to a sensitive word list. Therefore, the goal of attackers is to construct an adversarial prompt $\mathbf{p}_{\text{adv}}$ that does not contain any sensitive word  [^1] Attackers may incorporate any specific words into their sensitive word list during an attack, enabling them to effectively mask their malicious intentions., while leading the generation toward the semantics of the target prompt.
@@ 3.3 Text-Modal Attack
Semantic similarity-driven loss. We begin by inputting a target prompt $\mathbf{p}_{\text{tar}}$ that describes the desired content from the attacker’s perspective, as illustrated in Figure 3. To precisely reflect the attacker’s intentions, we formulate a targeted attack and utilize cosine similarity to ensure semantic similarity between $\mathbf{p}_{\text{adv}}$ and $\mathbf{p}_{\text{tar}}$. Our textual attack objective is formalized as:
$\vspace{-10pt}\max\cos(\tau_{\theta}(\mathbf{p}_{\text{tar}}),\tau_{\theta}(\mathbf{p}_{\text{adv}}))$ (1)
@@
Gradient-driven optimization. Inspired by the success of gradient-based adversarial attacks in computer vision (Madry et al., 2018; Carlini & Wagner, 2017; Goodfellow et al., 2015), it is important to utilize gradient information for effective attacks. However, the discrete nature of text tokens challenges the optimization of our defined objective in Equation 1. Inspired by gradient-driven optimization methods from the NLP domain like LLM-attack (Zou et al., 2023), FGPM (Wang et al., 2021), TextGrad (Hou et al., 2023), and prompt learning techniques (Shin et al., 2020), we harness token-level gradients to guide the optimization process. Specifically, we initiate the adversarial input sequence, $\mathbf{p}_{\text{adv}}=[p_{1},...,p_{i},...,p_{L}]$, with $L$ random tokens. For each token position $i$, every vocabulary token is considered as a potential substitute. A position-wise token selection variable, $\mathbf{s_{i}}=[s_{i1},s_{i2},...,s_{i|V|}]$ is introduced where $s_{ij}=1$ if the $j^{th}$ token is chosen at position $i$. We enable the gradient of all $\mathbf{s_{i}}$, and perform backpropagation on the attack loss function to calculate the gradient w.r.t $s_{ij}$ which is then used to measure the impact of the $j^{th}$ candidate token at position $i$. To search substitutional tokens, we utilize a greedy search strategy (Zou et al., 2023; Jin et al., 2020; Li et al., 2020; Garg & Ramakrishnan, 2020). Tokens are ranked by their gradients and the top $k$ tokens at each position are selected, creating a candidate prompt pool $\mathcal{P}$ of dimensions $\mathbb{N}^{L\times k}$. We then sample $q$ candidate prompts from $\mathcal{P}$, rank them according to their loss values, and choose the prompt $\mathbf{c}_{opt}$ with the highest optimization value in Equation 1. This prompt is set as $\mathbf{p}_{\text{adv}}$ for a single optimization iteration, and the process is repeated until the final adversarial prompt is obtained.
@@
Sensitive word regularization. To eliminate sensitive words in $\mathbf{p}_{\text{adv}}$, we construct a list of sensitive words based on the NSFW concepts investigated by  (Rando et al., 2022b; Qu et al., 2023), which typically includes explicit NSFW words, as highlighted in red font in  Figure 3 (see Appendix for the full word list). Later, we suppress the occurrence of tokens from the sensitive word list by setting their gradients to $-\inf$. As will be evident later, this sensitive words elimination strategy can effectively evade prompt filters, despite being implemented by advanced deep neural networks, as the AI moderator employed in Midjounery (Midjourney, 2023) and Leonardo.Ai (Leonardo.Ai, 2023).
@@ 3.4 Image-Modal Attack
T2I models like SD can use a post-hoc safety checker to identify NSFW content in the synthesis, replacing flagged synthesis with a black image as in  Figure 2 (b). This defense mechanism in image space motivates us to initiate attacks on the image modality to cheat these safety checkers.
In this image-modal attack, our focus is the image editing task of T2I models. Given that the image is prone to NSFW contents induced by malicious prompts, we aim to evade the post-hoc safety checker through the adversarial attack. As illustrated in Figure 4, given an NSFW-related prompt $\mathbf{p}$ and an input image $\mathbf{x}_{\text{input}}$, a T2I model generates a synthetic image, $\mathbf{x}_{\text{syn}}$. The safety checker then maps this image to a latent vector $I$ and compares it with $M$ default NSFW embeddings, denoted as $C_{i}$ for $i=1,...,M$, via cosine distance. If any cosine value exceeds the corresponding threshold $T_{i}$, the synthesis is flagged as NSFW. We expect the victim safety checker to release the synthesis $\mathbf{x}_{\text{syn}}$ by crafting $\mathbf{x}_{\text{adv}}$ as the model input. To achieve this objective, we dynamically optimize the gradients of loss items that exceed $T_{i}$, as shown in the red box in  Figure 4. We formulate our objective in Equation 2.
$\small\underset{\left\|\textbf{x}_{\text{input}}-\textbf{x}_{\text{adv}}\right\|_{2}\leq\varepsilon}{\arg\min}\sum_{i=1}^{M}\textbf{1}_{\left\{\cos(I,C_{i})>T_{i}\right\}}\cos(I,C_{i}),\vspace{-8pt}$ (2)
where 1 is the indicator function to select the triggered loss items for optimization, $\varepsilon$ indicates the perturbation budget. This dynamic loss selection strategy focuses on optimizing features near the decision boundary, allowing us to bypass the safety checker while minimally altering image features. The constrained optimization problem in  Equation 2 is solved using projected gradient descent (Madry et al., 2018) (detailed algorithm is provided in Appendix).
[…]
Metric Q16 Mhsc SC Avg.

### v4

@@ Abstract
In recent years, Text-to-Image (T2I) models have seen remarkable advancements, gaining widespread adoption. However, this progress has inadvertently opened avenues for potential misuse, particularly in generating inappropriate or Not-Safe-For-Work (NSFW) content. Our work introduces MMA-Diffusion, a framework that presents a significant and realistic threat to the security of T2I models by effectively circumventing current defensive measures in both open-source models and commercial online services. Unlike previous approaches, MMA-Diffusion leverages both textual and visual modalities to bypass safeguards like prompt filters and post-hoc safety checkers, thus exposing and highlighting the vulnerabilities in existing defense mechanisms. Our codes are available at https://github.com/cure-lab/MMA-Diffusion.
@@ 1 Introduction
In the rapidly evolving landscape of text-to-image (T2I) generation, diffusion models such as Stable Diffusion (SD) [30] and Midjounery [1] have marked a paradigm shift. These models have revolutionized digital creativity by generating strikingly realistic images, yet they also pose significant security challenges. Notably, the potential misuse of these models for generating Not-Safe-For-Work (NSFW) contents [34, 25, 31], such as adult materials, violence, and politically sensitive imagery, is a serious concern.
In response to these concerns, developers of T2I models have implemented preventive measures like prompt filters and post-synthesis safety checks. While effective to an extent, the resilience of these measures against sophisticated adversarial attacks remains a topic of intense debate and investigation. Our study delves into this pressing issue by introducing MMA-Diffusion, a framework designed to rigorously test and challenge the security of T2I models. Unlike conventional methods that make subtle prompt modifications [19, 8, 41, 14], MMA-Diffusion adopts a systematic attack approach. It enables users to generate unrestricted adversarial prompts and craft image perturbations, thereby circumventing existing safety protocols.
The technical prowess of MMA-Diffusion lies in its dual-modal attack strategy. We develop an advanced text modality attack mechanism that intricately alters textual prompts while maintaining their semantic intent, allowing for the generation of targeted NSFW content without being flagged by existing filters, as demonstrated in  Fig. 1(a). On the image modality front, MMA-Diffusion utilizes a novel perturbation technique that subtly alters image characteristics in a manner undetectable to the human eye but significant enough to bypass post-processing safety algorithms, as illustrated in  Fig. 1(b).
@@
Overall, the contributions of this work include: 1 We present a novel multimodal systematic attack that effectively bypasses prompt filters and safety checkers, highlighting a significant security issue in T2I models. 2 In the textual modality, we craft an adversarial prompt generation method that can deceive the prompt filter while remaining semantically similar to the target. For the image modality, we devise an attack that proficiently bypasses the post-hoc defense mechanism. 3 We evaluate various T2I models, encompassing popular open-source models and online platforms and demonstrate the effectiveness of the proposed MMA-Diffusion. For example, 10-query black-box attack can achieve a 83.33% and 90% success rate with respect to Midjounery [1] and Leonardo.Ai [3].
Figure 2: Overview of the proposed framework. T2I models incorporate safety mechanisms, including (a) prompt filters to prohibit unsafe prompts/words, e.g.naked," and (b) post-hoc safety checkers to prevent explicit synthesis. (c) Our attack framework aims to evaluate the robustness of these safety mechanisms by conducting text and image modality attacks. Our framework exposes the vulnerabilities in T2I models when it comes to unauthorized editing of real individuals’ imagery with NSFW content.
@@ 2 Related Work
Adversarial attacks on T2I models. To the best of our knowledge, current research does not extensively explore attacks in the image modality for NSFW content generation with T2I models. Most existing studies on adversarial attacks in T2I models, such as [32, 22, 25, 8, 41, 14, 20, 18], have predominantly focused on text modification to probe functional vulnerabilities. These explorations encompass impacts from diminishing synthetic quality [20, 39, 32] to distorting or eliminating objects [41, 20, 22], and impairing image fidelity [19, 22, 18]. However, they do not target generating NSFW-specific materials like pornography, violence, politics, racism, or horror. Recent works such as UnlearnDiff [40] and Ring-A-Bell  [37] have started to consider the misuse of T2I models for generating NSFW content. UnlearnDiff primarily examines concept-erased diffusion models [30, 7, 15, 34] and does not extend to other defense strategies. Conversely, Ring-A-Bell [37] explores inducing T2I models to generate NSFW concepts but lacks precision in controlling the details of the synthesis. However, none of them considers attacks that can bypass both the prompt filter and the post-hoc safety mechanisms while still producing high-quality NSFW content tailored to specific semantic prompts. This paper demonstrates the feasibility of such attacks, highlighting their general applicability across a variety of T2I models.
@@
Defensive methods. Various T2I models implement distinct countermeasures to mitigate user abuse. Notably, popular online T2I services like Midjourney [1] and Leonardo.Ai [3] employ AI moderators to screen potentially harmful prompts. This proactive approach targets the prevention of NSFW content generation at the input stage. Another defensive strategy involves post-hoc safety checkers, exemplified by the one integrated into Stable Diffusion (SD) [4, 28]. Unlike AI moderators, these checkers function at the output stage, scrutinizing generated images to detect and obfuscate NSFW elements. Additionally, some novel mitigation methods lie in the concept-erased diffusion [7, 15, 34]. These methods differ fundamentally from external safety measures as they modify the model’s inference guidance or utilize fine-tuning to actively suppress NSFW concepts. However, they may not entirely eliminate NSFW content and could inadvertently affect the quality of benign images [40, 16, 34]. This paper presents a multimodal attack that breaches both prompt filters and post-hoc safety checkers, which is also applicable to concept-erased diffusion models (e.g., SLD [34]), exposing the risk of T2I models and related online services.
@@ 3.1 Threat Model
  • White-Box Settings: Here, adversaries utilize open-source T2I models like SDv1.5 [5] for image generation. With full access to the model’s architecture and checkpoint, attackers can conduct in-depth explorations and manipulations for sophisticated attacks.
@@ 3.2 Approach Overview
In this paper, we focus on the attack that enables T2I models to generate high-quality NSFW content, thereby exposing the potential misuse risks of them, as in Fig. 2. Specifically, we assume that the attacker describes the content they wish to generate through plain text. The attack is considered successful only if the model generates NSFW content that aligns with the description.
To make the attack more realistic, we assume that the T2I model or the online service adopts two defense methods, namely: prompt filter, as in Fig. 2 (a) and post-hoc safety checker, as in Fig. 2 (b). For situations where only the prompt filter is present, such as [3], we employ a text-modal attack. For situations where only the post-hoc safety checker is present, such as SD [30], we utilize an image-modal attack. For models that adopt both modalities of defense, we can simultaneously use both attack methods to achieve a stronger effect, as in Fig. 2 (c).
3.3 Text-Modal Attack
T2I models typically rely on a pre-trained text encoder, $\tau_{\theta}(\cdot)$, to transforms natural language input $\mathbf{p}$ into a latent vector, denoted as $\tau_{\theta}(\mathbf{p})\in\mathbb{R}^{d}$, which is responsible for determining the semantics of the image synthesis [23]. The input sequence is $\mathbf{p}=[p_{1},p_{2},...,p_{L}]\in\mathbb{N}^{L}$, where $p_{i}\in\{0,1,...,|V|-1\}$ is the $i^{\text{th}}$ token’s index, $V$ is the vocabulary codebook, $|V|$ is the vocabulary size, and $L$ is the prompt length. This mapping from $\mathbb{N}^{L}$ to $\mathbb{R}^{d}$ provides a large search space for an attack, given a sufficiently large vocabulary pool $V$ and no additional constraints on $\mathbf{p}_{\text{adv}}$, thus enabling free-style adversarial prompt manipulation.
The target of the text-modal attack is to evade the prompt filter while keeping the functionality guiding the T2I model for the desired NSFW content. Specifically, we set this original NSFW prompt as the target prompt, denoted as $\mathbf{p}_{\text{tar}}$ (e.g., “A completely naked Trump stands on the grass"). MMA-Diffusion assumes the prompt filter is implemented by filtering the prompts according to a sensitive word list. Therefore, the goal of attackers is to construct an adversarial prompt $\mathbf{p}_{\text{adv}}$ that does not contain any sensitive word  [^1] Attackers may incorporate any specific words into their sensitive word list during an attack, enabling them to effectively mask their malicious intentions., while leading the generation toward the semantics of the target prompt.
@@ 3.3 Text-Modal Attack
Semantic similarity-driven loss. We begin by inputting a target prompt $\mathbf{p}_{\text{tar}}$ that describes the desired content from the attacker’s perspective, as illustrated in Fig. 3. To precisely reflect the attacker’s intentions, we formulate a targeted attack and utilize cosine similarity to ensure semantic similarity between $\mathbf{p}_{\text{adv}}$ and $\mathbf{p}_{\text{tar}}$. Our textual attack objective is formalized as:
$\max\cos(\tau_{\theta}(\mathbf{p}_{\text{tar}}),\tau_{\theta}(\mathbf{p}_{\text{adv}}))$ (1)
@@
Gradient-driven optimization. Inspired by the success of gradient-based adversarial attacks in computer vision [21, 6, 10], it is important to utilize gradient information for effective attacks. However, the discrete nature of text tokens challenges the optimization of our defined objective in Eq. 1. Inspired by gradient-driven optimization methods from the NLP domain like LLM-attack [42], FGPM [38], TextGrad [11], and prompt learning techniques [36], we harness token-level gradients to guide the optimization process. Specifically, we initiate the adversarial input sequence, $\mathbf{p}_{\text{adv}}=[p_{1},...,p_{i},...,p_{L}]$, with $L$ random tokens. For each token position $i$, every vocabulary token is considered as a potential substitute. A position-wise token selection variable, $\mathbf{s_{i}}=[s_{i1},s_{i2},...,s_{i|V|}]$ is introduced where $s_{ij}=1$ if the $j^{th}$ token is chosen at position $i$. We enable the gradient of all $\mathbf{s_{i}}$, and perform backpropagation on the objective to calculate the gradient w.r.t $s_{ij}$ which is then used to measure the impact of the $j^{th}$ candidate token at position $i$. To search substitutional tokens, we utilize a greedy search strategy [42, 12, 17, 9]. Tokens are ranked by their gradients and the top $k$ tokens at each position are selected, creating a candidate prompt pool $\mathcal{P}$ of $\mathbb{N}^{L\times k}$. We then sample $q$ candidate prompts from $\mathcal{P}$, rank them according to their loss values, and choose the prompt $\mathbf{c}_{opt}$ with the highest optimization value in Eq. 1. This prompt is set as $\mathbf{p}_{\text{adv}}$ for a single optimization iteration, and the process is repeated until the final adversarial prompt is obtained.
@@
Sensitive word regularization. To eliminate sensitive words in $\mathbf{p}_{\text{adv}}$, we construct a list of sensitive words based on the NSFW concepts investigated by  [29, 25], which typically includes explicit NSFW words, as highlighted in bold red font in  Fig. 3 (see Appendix for the full word list). Later, we suppress the occurrence of tokens from the sensitive word list by setting their gradients to $-\inf$. As will be evident later, this sensitive words elimination strategy can effectively evade prompt filters, despite being implemented by advanced deep neural networks, as the AI moderator employed in Midjounery [1] and Leonardo.Ai [3].
@@ 3.4 Image-Modal Attack
T2I models like SD can use a post-hoc safety checker to identify NSFW content in the synthesis, replacing flagged synthesis with a black image as in  Fig. 2 (b). This defense mechanism in image space motivates us to initiate attacks on the image modality to cheat these safety checkers.
In this image-modal attack, our focus is the image editing task of T2I models. Given that the image is prone to NSFW contents induced by malicious prompts, we aim to evade the post-hoc safety checker through the adversarial attack. As illustrated in Fig. 4, given an NSFW-related prompt $\mathbf{p}$ and an input image $\mathbf{x}_{input}$, a T2I model generates a synthetic image, $\mathbf{x}_{syn}$. The safety checker then maps this image to a latent vector $I$ and compares it with $M$ default NSFW embeddings, denoted as $C_{i}$ for $i=1,...,M$, via cosine distance. If any cosine value exceeds the corresponding threshold $T_{i}$, the synthesis is flagged as NSFW. We expect the victim safety checker to release the synthesis $\mathbf{x}_{syn}$ by crafting $\mathbf{x}_{adv}$ as the model input. To achieve this objective, we dynamically optimize the gradients of loss items that exceed $T_{i}$, as shown in the red box in  Fig. 4. We formulate our objective in Eq. 2.
$\small\underset{\left\|\textbf{x}_{\text{input}}-\textbf{x}_{\text{adv}}\right\|_{2}\leq\varepsilon}{\arg\min}\sum_{i=1}^{M}\textbf{1}_{\left\{\cos(I,C_{i})>T_{i}\right\}}\cos(I,C_{i}),$ (2)
where 1 is the indicator function to select the triggered loss items for optimization, $\varepsilon$ indicates the perturbation budget. This dynamic loss selection strategy focuses on optimizing features near the decision boundary, allowing us to bypass the safety checker while minimally altering image features. The constrained optimization problem in  Eq. 2 is solved using projected gradient descent [21]. Detailed algorithm is provided in Algorithm 1.
Algorithm 1 Image-modal Adversarial Attack
Input image $\mathbf{x}_{input}$, prompt $\mathbf{p}$, Stabel Diffusion model $SD$, CLIP’s vision encoder $\mathcal{V}_{en}$, predefined pornographic concept embeddings $C_{i},i=1,...,M$, predefined NSFW thresholds $T_{i},i=1,...,M$, perturbation budget $\varepsilon$, step size $\alpha$, number of iteration $N$.
Initialize $\textbf{x}_{adv}=\textbf{x}_{input}$
for $i=1,...,N$ do
    Generating the synthesis: $\mathbf{x}_{syn}\leftarrow SD(\mathbf{x}_{adv},\mathbf{p})$
    Obtain image embedding: $I\leftarrow\mathcal{V}_{en}(\textbf{x}_{syn})$
    Calculate loss: $\mathcal{L}\leftarrow\sum_{i=1}^{M}\textbf{1}_{\left\{\cos(I,C_{i})>T_{i}\right\}}\cos(I,C_{i})$
    Updating gradient: $\delta\leftarrow\delta+\alpha\cdot sign(\nabla_{\mathbf{x}_{adv}}\mathcal{L})$
    Projecting gradient: $\delta\leftarrow\text{clamp}(\delta,-\varepsilon,\varepsilon)$
    Update adversarial image: $\mathbf{x}_{adv}\leftarrow\mathbf{x}_{adv}-\delta$
end for
Optimized adversarial image $\mathbf{x}_{adv}$
4 Experiments
Metric Q16 [33] Mhsc [25] SC [4] Avg.